Tori bought some beach towels on sale for $8 each. Theo bought the same number of beach towels at the full price of $12. Toni's total was $24 less than Theo's total. How many beach towels did they each buy?
Answer by ikleyn(52810)

.
Let n be the number of towels, same for both persons.


Tori spent 8n dollars.  Theo spent 12n dollars.


The difference is 24 dollars:


    12n - 8n = 24.


    4n = 24

     n = 24/4 = 6.     


ANSWER.  6 towels.